# SPC: Evolving Self-Play Critic via Adversarial Games for LLM Reasoning

### v1

@@ Abstract
Evaluating the step-by-step reliability of large language model (LLM) reasoning, such as Chain-of-Thought, remains challenging due to the difficulty and cost of obtaining high-quality step-level supervision. In this paper, we introduce Self-Play Critic (SPC), a novel approach where a critic model evolves its ability to assess reasoning steps through adversarial self-play games, eliminating the need for manual step-level annotation. SPC involves fine-tuning two copies of a base model to play two roles, namely a “sneaky generator” that deliberately produces erroneous steps designed to be difficult to detect, and a “critic” that analyzes the correctness of reasoning steps. These two models engage in an adversarial game in which the generator aims to fool the critic, while the critic model seeks to identify the generator’s errors. Using reinforcement learning based on the game outcomes, the models iteratively improve; the winner of each confrontation receives a positive reward and the loser receives a negative reward, driving continuous self-evolution. Experiments on three reasoning process benchmarks (ProcessBench, PRM800K, DeltaBench) demonstrate that our SPC progressively enhances its error detection capabilities (e.g., accuracy increases from 70.8% to 77.7% on ProcessBench) and surpasses strong baselines, including distilled R1 model. Furthermore, applying SPC to guide the test-time search of diverse LLMs significantly improves their mathematical reasoning performance on MATH500 and AIME2024, outperforming state-of-the-art process reward models.
@@ 1 Introduction
The Chain-of-Thought (CoT) [1, 2, 3] reasoning process, which emerges in the autoregressive generation of large language models (LLMs), has been applied to address a variety of complex tasks [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Training methods such as Supervised Fine-Tuning (SFT) [14, 15], Reinforcement Learning from Human Feedback (RLHF) [16, 17], and self-play reinforcement learning [18, 19], have demonstrated success in obtaining high-quality CoT. Recently, the popular o1 [20], R1 [21], and QwQ [22] LLMs utilize large-scale reinforcement learning for training and employ test-time scaling to generate long CoT, further enhancing their reasoning capabilities. As the CoT generated by LLMs becomes increasingly complex and diverse, it is particularly important to verify the reliability of the reasoning process, analyze the potential errors in reasoning steps, and guide the test-time search to improve the reasoning process [23, 24, 25, 26, 27, 28, 29, 30, 31].
A number of verification models have been developed to analyze and evaluate the reasoning process of LLMs. For example, outcome verifiers [25] provide outcome-level validation to rank or reward multiple responses from LLMs. Process verifiers [25, 23], which validate each step in the reasoning process, have proven crucial in recent advances in LLM reasoning [26, 32, 33, 34]. However, there are several challenges that limit the development of such step-level approaches. Firstly, while it is relatively simple to extract the final answer to determine the correctness of a solution and automatically collect training data, determining the correctness of a reasoning step and obtaining well-annotated step data for training a process verifier is much more difficult. Secondly, LLMs are updated rapidly, and heavy human expert annotations on the outputs of specific LLMs may not be applicable to the latest LLMs due to distributional differences. Thirdly, the dataset limited to step correctness annotations restricts the training of a critic model – preventing it from providing substantive feedback and reducing it to merely a scoring mechanism for verification.
@@ 2 Related Work
Powerful large language models (LLMs) [4, 5, 7, 8, 9, 10, 11, 12, 13] not only excel at following human instructions to handle various general tasks, but they are also becoming increasingly adept at constructing Chain-of-Thought (CoT) to tackle complex reasoning tasks, such as solving math problems and code generation. The recently popular o1 [20], R1 [21], and QwQ [22] models are further equipped with exceptional deep thinking capabilities, allowing them to construct long CoT during inference to decompose complex tasks and even perform extensive self-critique and self-correction. However, fine-grained analyses in a recent research [35] indicate that the effective proportion of self-critique in these long CoT is still very low, and biases exist in the self-critique of their own reasoning processes. It is therefore necessary to have a simple external critic for assessing the reasoning steps of various LLMs, providing step-level critiques.
@@
Verifiers [23, 24, 25, 26, 27] can enhance reasoning performance by scoring the reasoning process, thereby allowing for the ranking or integration of multiple responses generated by LLMs during inference. Additionally, they can also provide more accurate rewards during training to guide the optimization of the LLM. Verifiers can be primarily categorized into two types, namely outcome reward models (ORMs) and process reward models (PRMs). ORMs provide solution-level scores for the entire problem-solving process, whereas PRMs assign step-level scores to each step of the reasoning process, which can be aggregated to produce a more accurate solution-level score. Recent works [28, 29, 39, 30] propose critic models for verification, arguing that scalar scores have limited ability in evaluating the outputs of LLMs. In contrast, feedback in natural language form can activate the thinking capabilities of LLMs, resulting in more reliable critiques to represent the correctness of reasoning. In this work, we explore how to analyze the correctness of the current step based on partial reasoning steps. This allows us to acquire step-level critiques to help improve the performance of LLMs during test-time search.
@@
Self-play [40, 41] is a method in reinforcement learning where an agent interacts with several copies of itself in an environment to learn specific actions. A significant advancement in self-play is demonstrated by AlphaGo [42] and AlphaZero [43], which greatly surpass human champions in the game of Go, without the need for human knowledge in training. Recent studies apply self-play to LLM alignment and enhancement [19, 18, 44, 45, 46]. For example, Kirchner et al. [19] proposed a solution-level game between a powerful generator and a weak scoring verifier to enhance the legibility of the LLM, though resulting in performance degradation. Cheng et al. [44] introduced a Taboo language game between an attacker and a defender to improve LLMs’ reasoning abilities. In this paper, we design an adversarial game to generate data for training a step-level critic, which provides correctness analysis for the reasoning steps of LLMs.
@@ 3.1 Overview
Sneaky generator converts correct reasoning steps from LLMs into incorrect ones, automating the creation of numerous steps with potential errors. Its goal is to generate sneaky steps that not only decrease the reasoning success rate of LLMs but also deceive the critic (i.e., the critic fails to detect the erroneous steps). Specifically, given a problem $p$ and a correct partial reasoning trajectory $\tau_{:k}=(t_{1},t_{2},...t_{k})$ produced by an LLM solver, the sneaky generator $S$ converts the last correct step $t_{k}^{c}$ into a candidate sneaky step $t_{k}^{i}=S(p,\tau_{:k-1},t_{k}^{c})$. This candidate becomes a valid sneaky step $\overline{t_{k}^{i}}$ if it significantly impacts the solver’s success rate in the subsequent competition from this step. If the sneaky step is invalid or the critic detects the errors in a valid sneaky step, the sneaky generator receives a negative reward. Conversely, if the critic fails to detect the errors in a valid sneaky step, the sneaky generator receives a positive reward. By automatically evaluating the success of the generated sneaky steps, we then employ reinforcement learning to enable the self-evolution of the sneaky generator.
@@ 3.2 Initializing Sneaky Generator
To initialize the sneaky generator $S_{0}$, we train the base model Qwen2.5-7B-Instruct [12] using Supervised Fine-Tuning (SFT) to equip it with the fundamental capability to generate incorrect steps. To ensure the accuracy of the initialization data, we use correct-incorrect step pairs from PRM800K [23] to construct an error step transformation process. Specifically, we extract correct-incorrect step pairs $<t_{k}^{c},{t_{k}^{i}}>$ with the same problem $p$ and partial solution $\tau_{:k-1}=(t_{1},t_{2},...t_{k-1})$ (the steps preceding the extracted pairs) from PRM800K. We next prompt GPT-4 to create a chain-of-thought transformation $\mathcal{T}_{\text{CoT}}(t_{k}^{c})\rightarrow{t_{k}^{i}}$ by first selecting an error type from five predefined common error types (see Sec. A) and then performing a detailed transformation. This process results in a transformation behavior cloning dataset $(\mathbf{x},\mathbf{y})\sim\mathcal{D}_{bc}^{S}$, where $\mathbf{x}=(p,\tau_{:k-1},t_{k}^{c})$ as input, $\mathbf{y}=\mathcal{T}_{\mathrm{CoT}}(t_{k}^{c})\rightarrow{t_{k}^{i}}$ as output. We then finetune Qwen2.5-7B-Instruct on dataset $\mathcal{D}_{bc}^{S}$ to obtain a policy $\pi_{\theta}$ for the initial sneaky generator $S_{0}$ using the SFT loss:
@@
To form an adversarial game, we need to annotate the generated steps and feed actual incorrect steps to the critic model. However, existing LLM-as-a-Judge methods [47, 35] inevitably introduce bias, while the manual annotation is excessively labor-intensive. We therefore propose evaluating the impact of different steps on the problem-solving success rate to ascertain whether a sneaky step can be considered incorrect. Concretely, based on a correct solution generated by an open-source LLM, we first sample an original step and transform it into a sneaky step using the sneaky generator. We subsequently use the same LLM to complete the entire reasoning process after the original/sneaky steps, and this is repeated $N$ times. If the original step achieves a relatively high success rate while the sneaky step results in a significantly lower success rate, we consider this pair of steps to represent correct and incorrect steps, respectively. In our experiment, we adopted a strict criterion to ensure data quality. If the original step achieves a success rate greater than or equal to 75%, while the sneaky step results in a success rate of 0%, we then collect this pair of steps for subsequent adversarial games.
@@ 3.3 Initializing Step Critic
Specifically, we prompt DeepSeek-R1-Distill-Qwen-7B as a critic, taking problem $p$, partial solutions $\tau_{:k-1}$, and mixed correct and incorrect steps $t_{k}$ from PRM800K dataset as inputs, to collect long critiques. We then employ GPT-4 to rewrite them into brief standardized critiques $Q_{t}$ (see Sec. A). For example, the first part of a standardized critique is an analysis of the partial solution, the second part provides an analysis of the current last step, and the third part gives a definite conclusion regarding the correctness of the step. This also simplifies the task and facilitates the use of SFT for policy initialization. Additionally, when preparing the training data for the critic, we mix the steps labeled as correct and incorrect in PRM800K at a 1:1 ratio to ensure the critic’s capabilities are balanced. We utilize human annotations from PRM800K to filter around 21.8K correctly generated critiques $\overline{Q_{t}}$. Similarly, we prepare behavior cloning dataset $(\mathbf{x},\mathbf{y})\sim\mathcal{D}_{bc}^{C}$ for the critic, where $\mathbf{x}=(p,\tau_{:k-1},t_{k})$ as input, and $\mathbf{y}=\overline{Q_{t}}$ as output. We then finetune the base model using SFT loss (1) to obtain an initial policy $C_{0}$ for the critic.
@@ 3.4 Adversarial Game
$R_{critic}$ =$=\begin{cases}1,&\texttt{Critic Wins}\\ -1,&\texttt{Critic Loses}\\ \end{cases}$ (3)
@@ 3.5 Evolving via Reinforcement Learning
In each iteration, after obtaining positive and negative samples through the adversarial games, we apply offline reinforcement learning to the critic and sneaky generator, respectively, enabling self-improvement of both roles based on the game result. Specifically, we adopt the following optimization objective to achieve efficient and stable RL training:
@@
where $\pi_{\text{old}}$ denotes the policy used to collect the offline dataset, $\frac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\text{old}}(\mathbf{y}|\mathbf{x})}$ is the importance ratio, and $\hat{A}^{\pi_{\text{old}}}$ represents the advantage estimation  [48]. In practice, we adopt $\hat{A}^{\pi_{\text{old}}}=R(\mathbf{x},\mathbf{y})-\beta\text{KL}[\pi_{\theta}\|\pi_{\text{ref}}]$ for the advantage estimation, where a Kullback-Leibler (KL) penalty is added to regularize the policy $\pi_{\theta}$ and prevent it from deviating too far from the initial policy $\pi_{\text{ref}}$ [49].
@@
Based on the adversarial game, we apply iterative training to enable continuous evolution of the two roles. Specifically, in each iteration, the newly updated policies engage in the adversarial game to further generate data, updating the initial policy into an evolved version. Additionally, we observe an interesting phenomenon that more balanced adversarial games contribute to the self-evolution of models. In fact, the initial sneaky generator $S_{0}$ is weaker than the initial critic $C_{0}$, resulting in an unbalanced win rate. Moreover, $S_{1}$ obtained through synchronous iteration is even weaker than $C_{1}$. Therefore, we adopt an asymmetric evolution strategy, where the stronger $S_{1}$ competes against $C_{0}$ in a more balanced game to generate the second round of data. This enables the critic $C_{2}$ trained in the second round to further improve its performance. Such a strategy is analogous to humans preferring to improve their skills in chess by playing against equally matched opponents. We provide more detailed analyses of adversarial games in Sec. 4.3.
@@ 4.1 Experimental Settings
Evaluation   We adopt PRM800K [23], ProcessBench [27], and DeltaBench [35] that include human annotations of mathematical reasoning steps for evaluation. The original setting of ProcessBench and DeltaBench is to identify the position of the first or all errors in a complete solution. We argue that, in practical scenarios, a critic can enhance reasoning performance by identifying the incorrect step and requiring the LLM to regenerate, with no need to wait for completing all the steps. We therefore extract a 1:1 ratio of correct and erroneous steps from each benchmark, only retain the reasoning process before these steps as a partial solution, and discard the reasoning steps after these steps. Besides, we evaluate the effectiveness of the critic models in assisting LLMs to solve math problems on MATH500 [50] and AIME2024 [37]. More evaluation details are provided in Sec. B.
Baselines   Following ProcessBench [27], we primarily evaluate two types of baselines, namely Process Reward Models (PRMs) and prompting LLMs as critic models. For PRMs, we select two representative methods, namely Math-Shepherd [26] and Qwen2.5-Math-7B-PRM800K [27]. Math-Shepherd trains a process reward model through an automated data annotation process and can be utilized to rank multiple outputs or ensemble them to enhance reasoning performance. Qwen2.5-Math-7B-PRM800K is based on the advanced math-specialized model Qwen2.5-Math-7B [51], and is further fine-tuned with the PRM800K dataset, obtaining state-of-the-art performance among PRMs. We also prompt multiple types of LLMs to serve as critic models, using the same prompts in our critic models’ training. Several representative models, including Llama [10], Qwen [12], R1 [21], and GPT-4o [6], are selected as baselines.
Table 1: Comparison of recall on ProcessBench. We evaluate different models on their ability to assess the correctness of the current step during reasoning search, instead of only predicting the index of the first error in the complete solution. ‘Round 0’ refers to the initialized critic model.

### v2

@@ Abstract
Evaluating the step-by-step reliability of large language model (LLM) reasoning, such as Chain-of-Thought, remains challenging due to the difficulty and cost of obtaining high-quality step-level supervision. In this paper, we introduce Self-Play Critic (SPC), a novel approach where a critic model evolves its ability to assess reasoning steps through adversarial self-play games, eliminating the need for manual step-level annotation. SPC involves fine-tuning two copies of a base model to play two roles, namely a “sneaky generator” that deliberately produces erroneous steps designed to be difficult to detect, and a “critic” that analyzes the correctness of reasoning steps. These two models engage in an adversarial game in which the generator aims to fool the critic, while the critic model seeks to identify the generator’s errors. Using reinforcement learning based on the game outcomes, the models iteratively improve; the winner of each confrontation receives a positive reward and the loser receives a negative reward, driving continuous self-evolution. Experiments on three reasoning process benchmarks (ProcessBench, PRM800K, DeltaBench) demonstrate that our SPC progressively enhances its error detection capabilities (e.g., accuracy increases from 70.8% to 77.7% on ProcessBench) and surpasses strong baselines, including distilled R1 model. Furthermore, SPC can guide the test-time search of diverse LLMs and significantly improve their mathematical reasoning performance on MATH500 and AIME2024, surpassing those guided by state-of-the-art process reward models.
@@ 1 Introduction
The Chain-of-Thought (CoT) [1, 2, 3] reasoning process, which emerges in the autoregressive generation of large language models (LLMs), has been applied to address a variety of complex tasks [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Training methods such as Supervised Fine-Tuning (SFT) [14, 15], Reinforcement Learning from Human Feedback (RLHF) [16, 17], and self-play reinforcement learning [18, 19], have demonstrated success in obtaining high-quality CoT. Recently, the popular o1 [20], R1 [21], and QwQ [22] utilize large-scale reinforcement learning for training and employ test-time scaling to generate long CoT, further enhancing their reasoning capabilities. As the CoT generated by LLMs becomes increasingly complex and diverse, it is particularly important to verify the reliability of the reasoning process, analyze the potential errors in reasoning steps, and guide the test-time search to improve the reasoning process [23, 24, 25, 26, 27, 28, 29, 30, 31].
A number of verification models have been developed to analyze and evaluate the reasoning process of LLMs. For example, outcome verifiers [25] provide outcome-level validation to rank or reward multiple responses from LLMs. Process verifiers [25, 23], which validate each step in the reasoning process, have proven crucial in recent advances in LLM reasoning [26, 32, 33, 34]. However, there are several challenges that limit the development of such step-level approaches. Firstly, while it is relatively simple to extract the final predicted answer and determine the correctness of a solution, determining the correctness of a reasoning step and automatically obtaining well-annotated step data for training a process verifier is much more difficult. Secondly, LLMs are updated rapidly, and heavy human expert annotations on the outputs of specific LLMs may not be applicable to the latest LLMs due to distributional differences. Thirdly, the dataset limited to step correctness annotations restricts the training of a critic model – preventing it from providing substantive feedback and reducing it to merely a scoring mechanism for verification.
@@ 2 Related Work
Powerful large language models (LLMs) [4, 5, 7, 8, 9, 10, 11, 12, 13] are becoming increasingly adept at constructing Chain-of-Thought (CoT) to tackle complex reasoning tasks, such as solving math problems and code generation. The recently popular o1 [20], R1 [21], and QwQ [22] models are further equipped with exceptional deep thinking capabilities, allowing them to construct long CoT during inference to decompose complex tasks and even perform extensive self-critique and self-correction [39, 40, 41]. However, fine-grained analyses in a recent research [35] indicate that the effective proportion of self-critique in these long CoT is still very low, and biases exist in the self-critique of their own reasoning processes. It is therefore necessary to have a simple external critic for assessing the reasoning steps of various LLMs, providing step-level critiques.
@@
Verifiers [23, 24, 25, 26, 27] can enhance reasoning performance by ranking or integrating multiple responses generated by LLMs during inference. Additionally, they can also provide more accurate rewards during training to guide the optimization of LLMs. Verifiers can be categorized into two types, namely outcome reward models (ORMs) and process reward models (PRMs). ORMs provide solution-level scores for the entire problem-solving process, whereas PRMs assign step-level scores to each reasoning step, which can be aggregated to produce a more accurate solution-level score. Recent works [28, 29, 42, 30] propose critic models for verification, arguing that scalar scores have limited ability in evaluating the outputs of LLMs. In contrast, feedback in natural language form can activate the thinking capabilities of LLMs, resulting in more reliable critiques to represent the correctness of reasoning. In this work, we explore how to analyze the correctness of the current step and provide step-level critiques based on partial reasoning steps.
@@
Self-play [43, 44] is a method in reinforcement learning where an agent interacts with several copies of itself in an environment to learn specific actions. A significant advancement in self-play is demonstrated by AlphaGo [45] and AlphaZero [46], which greatly surpass human champions in the game of Go, without the need for human knowledge in training. Recent studies apply self-play to LLM alignment and enhancement [19, 18, 47, 48, 49, 50]. For example, Kirchner et al. [19] proposed a solution-level game between a powerful prover and a weak scoring verifier to enhance the legibility of the LLM, though resulting in performance degradation. Cheng et al. [47] introduced a Taboo language game between an attacker and a defender to improve LLMs’ reasoning abilities. In this paper, we design an adversarial game to generate data for training a step-level critic, which provides correctness analysis for the reasoning steps of LLMs.
@@ 3.1 Overview
Sneaky generator converts correct reasoning steps from LLMs into incorrect ones, automating the creation of numerous steps with potential errors. Its goal is to generate sneaky steps that not only decrease the reasoning success rate of LLMs but also deceive the critic (i.e., the critic fails to detect the erroneous steps). Specifically, given a problem $p$ and a correct partial reasoning trajectory $\tau_{:k}=(t_{1},t_{2},...t_{k})$ produced by an LLM solver, the sneaky generator $S$ converts the last correct step $t_{k}^{c}$ into a candidate sneaky step $t_{k}^{i}=S(p,\tau_{:k-1},t_{k}^{c})$. This candidate becomes a valid sneaky step $\overline{t_{k}^{i}}$ if it significantly impacts the solver’s success rate in the subsequent completion from this step. If the sneaky step is invalid or the critic detects the errors in a valid sneaky step, the sneaky generator receives a negative reward. Conversely, if the critic fails to detect the errors in a valid sneaky step, the sneaky generator receives a positive reward. By automatically evaluating the success of the generated sneaky steps, we then employ reinforcement learning to enable the self-evolution of the sneaky generator.
@@ 3.2 Initializing Sneaky Generator
To initialize the sneaky generator $S_{0}$, we train the base model Qwen2.5-7B-Instruct [12] using Supervised Fine-Tuning (SFT) to equip it with the fundamental capability to generate incorrect steps. To ensure the accuracy of the initialization data, we use correct-incorrect step pairs from PRM800K [23] to construct an error step transformation process. Specifically, we extract correct-incorrect step pairs $<t_{k}^{c},{t_{k}^{i}}>$ with the same problem $p$ and partial solution $\tau_{:k-1}=(t_{1},t_{2},...t_{k-1})$ (the steps preceding the extracted pairs) from PRM800K. We next prompt GPT-4 to create a chain-of-thought transformation $\mathcal{T}_{\text{CoT}}(t_{k}^{c})\rightarrow{t_{k}^{i}}$ by first selecting an error type from five predefined common error types (see Sec. B) and then performing a detailed transformation. This process results in a transformation behavior cloning dataset $(\mathbf{x},\mathbf{y})\sim\mathcal{D}_{bc}^{S}$, where $\mathbf{x}=(p,\tau_{:k-1},t_{k}^{c})$ as input, $\mathbf{y}=\mathcal{T}_{\mathrm{CoT}}(t_{k}^{c})\rightarrow{t_{k}^{i}}$ as output. We then finetune Qwen2.5-7B-Instruct on dataset $\mathcal{D}_{bc}^{S}$ to obtain a policy $\pi_{\theta}$ for the initial sneaky generator $S_{0}$ using the SFT loss:
@@
To form an adversarial game, we need to annotate the generated steps and feed actual incorrect steps to the critic model. However, existing LLM-as-a-Judge methods [51, 35] inevitably introduce bias, while the manual annotation is excessively labor-intensive. We therefore propose evaluating the impact of different steps on the problem-solving success rate to ascertain whether a sneaky step can be considered incorrect. Concretely, based on a correct solution generated by an open-source LLM, we first sample an original step and transform it into a sneaky step using the sneaky generator. We subsequently use the same LLM to complete the entire reasoning process after the original/sneaky steps, and this is repeated $N$ times. If the original step achieves a relatively high success rate while the sneaky step results in a significantly lower success rate, we consider this pair of steps to represent correct and incorrect steps, respectively. In our experiment, we adopt a strict criterion to ensure data quality. If the original step achieves a success rate greater than or equal to 75%, while the sneaky step results in a success rate of 0%, we then collect this pair of steps for subsequent adversarial games.
@@ 3.3 Initializing Step Critic
Specifically, we prompt DeepSeek-R1-Distill-Qwen-7B as a critic, taking problem $p$, partial solutions $\tau_{:k-1}$, and mixed correct and incorrect steps $t_{k}$ from PRM800K dataset as inputs, to collect long critiques. We then employ GPT-4 to rewrite them into brief standardized critiques $Q_{t}$ (see Sec. B). Concretely, $Q_{t}$ contains an analysis of the partial solution and the current last step, as well as a definite conclusion regarding the correctness of the step. This also simplifies the task and facilitates the use of SFT for policy initialization. Additionally, when preparing the training data for the critic, we mix the steps labeled as correct and incorrect in PRM800K at a 1:1 ratio to ensure the critic’s capabilities are balanced. We utilize human annotations from PRM800K to filter around 21.8K correctly generated critiques $\overline{Q_{t}}$. Similarly, we prepare behavior cloning dataset $(\mathbf{x},\mathbf{y})\sim\mathcal{D}_{bc}^{C}$ for the critic, where $\mathbf{x}=(p,\tau_{:k-1},t_{k})$ as input, and $\mathbf{y}=\overline{Q_{t}}$ as output. We then finetune the base model using SFT loss (1) to obtain an initial policy $C_{0}$ for the critic.
@@ 3.4 Adversarial Game
$R_{critic}$ =$=\begin{cases}1,&\texttt{Critic Wins}\\ -1,&\texttt{Critic Loses}\end{cases}$ (3)
@@ 3.5 Evolving via Reinforcement Learning
In each iteration, after obtaining positive and negative samples through the adversarial games, we apply offline reinforcement learning to the critic and sneaky generator, respectively, enabling self-improvement of both roles based on the game result. Specifically, we adopt the following optimization objective to achieve stable RL training:
@@
where $\pi_{\text{old}}$ denotes the policy used to collect the offline dataset, $\frac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\text{old}}(\mathbf{y}|\mathbf{x})}$ is the importance ratio, and $\hat{A}^{\pi_{\text{old}}}$ represents the advantage estimation. Inspired by recent RLOO [52] and GRPO [53], we formulate $\hat{A}^{\pi_{\text{old}}}=R(\mathbf{x},\mathbf{y})-b-\beta\text{KL}[\pi_{\theta}\|\pi_{\text{ref}}]$, where a baseline $b$ (the average reward of all samples) is subtracted for advantage estimation, and a Kullback-Leibler (KL) penalty is added to regularize the policy $\pi_{\theta}$ and prevent it from deviating too far from the initial policy $\pi_{\text{ref}}$.
@@
Based on the adversarial game, we apply iterative training to enable continuous evolution of the two roles. Specifically, in each iteration, the newly updated policies re-engage in the adversarial game to generate new data for training, thereby evolving themselves further. Additionally, we observe an interesting phenomenon that more balanced adversarial games contribute to the self-evolution of models. In fact, the initial sneaky generator $S_{0}$ is weaker than the initial critic $C_{0}$, resulting in an unbalanced win rate. Moreover, $S_{1}$ obtained through synchronous iteration is even weaker than $C_{1}$. Therefore, we adopt an asymmetric evolution strategy, where $S_{1}$ competes against $C_{0}$ in a more balanced game to generate the second round of data. This enables $C_{2}$ trained in the second round to further improve its performance. Such a strategy is analogous to humans preferring to improve their skills in chess by playing against equally matched opponents. We provide more detailed analyses of the evolving strategies in Sec. 4.3.
@@ 4.1 Experimental Settings
Evaluation
We adopt PRM800K [23], ProcessBench [27], and DeltaBench [35] that include human annotations of mathematical reasoning steps for evaluation. The original setting of ProcessBench and DeltaBench is to identify the position of the first or all errors in a complete solution. We argue that, in practical scenarios, a critic can enhance reasoning performance by identifying the incorrect step and requiring the LLM to regenerate, with no need to wait for completing all the steps. We therefore extract a 1:1 ratio of correct and erroneous steps from each benchmark, only retain the reasoning process before these steps as a partial solution, and discard the reasoning steps after these steps. Besides, we evaluate the effectiveness of the critic models in assisting LLMs to solve math problems on MATH500 [54] and AIME2024 [37]. More evaluation details are provided in Sec. C.
Baselines
Following ProcessBench [27], we primarily evaluate two types of baselines, namely Process Reward Models (PRMs) and prompting LLMs as critic models. For PRMs, we select two representative methods, namely Math-Shepherd [26] and Qwen2.5-Math-7B-PRM800K [27]. Math-Shepherd trains a process reward model through an automated data annotation process and can be utilized to rank multiple outputs or ensemble them to enhance reasoning performance. Qwen2.5-Math-7B-PRM800K is based on the advanced math-specialized model Qwen2.5-Math-7B [55], and is further fine-tuned with the PRM800K dataset, obtaining state-of-the-art performance among PRMs. We also prompt multiple types of LLMs to serve as critic models, using the same prompts as our SPC. Several representative models, including Llama [10], Qwen [12], R1 [21], and GPT-4o [6], are selected as baselines.
Table 1: Comparison of recall on ProcessBench. We evaluate different models on their ability to assess the correctness of the current step, instead of only predicting the index of the first error in the complete solution. ‘Round 0’ refers to the initialized critic model.